if you have $3.45 In quarters and dimes, and he has four more quarters than dimes how much does he have of each coin?

Answer by robertb(5830)

q = d + 4,
0.25q+0.10d = 3.45.
By direct substitution, 0.25(d+4) + 0.10d = 3.45,
0.25d+1 + 0.10d = 3.45,
0.35d = 2.45,
d = 7, the number of dimes.
q = 11, the number of quarters.